Natalia and Yara argue over who sits in the front seat of the car with their father. To settle the argument their father suggests the following method. Natalia tosses a coin twice and Yara makes a call on the outcome. If Yara calls both tosses correctly. Yara sits in the front. If not, then Yara chooses a number from one to six and rolls a die. If Yara rolls chosen number, she sits in the front seat. If Yara does not roll that number, she chooses two colours. She then draws a disc from a bag that contains five different coloured discs. If she draws out one of the colours she chose, Yara sits in the front seat. If not, Yara sits in the back seat. What is the probability of Yara sitting in the front seat?
A. 1/4
B. 2/5
C. 1/2
D. 5/8

Answer by Boreal(15235)

Probability=(1/4)+[(1/6)*(3/4)]=(1/4)+(1/8)=(3/8). That is the probability after the dice.
Choosing two colors out of 5: (2/5)
Probability is increased by (5/8)(2/5)=(1/4)
She has a (5/8) probability; (1/4) from getting the coins right, (1/8) from the dice, and (1/4) from the discs.
D